OpenStudy (anonymous):

find the slope of the line through the pair of points. (6,-9) and (-3,5)

OpenStudy (zepp):

Use slope = delta y / delta x

OpenStudy (zepp):

Delta y = y2 - y1 Delta x = x2 - x1

OpenStudy (anonymous):

\[slope= y2-y1 \div x2-x1 = 5--9\div -3-6\]= 14/-9=-1.56

OpenStudy (zepp):

@jesuskim m=[(-3)-(6)]/[(5)-(-9)] = -9/14 You did dx/dy :P

OpenStudy (anonymous):

whoops my mistake. -14/9 is correct. I went and did dx/dy instead of dy/dx sorry! thanks for finding it

OpenStudy (anonymous):

the slope of the line is given by the equation: slope(m) = rise/run = (dy/dx) dy means change in the y values, and dx means the change in the x values. so since the points are given in (x,y) fashion, you can set the equation up as follows: m=[(5)-(-9)]/[(-3)-(6)] = -14/9 goodluck!